Find the EAR in each of the following cases: (Assume 365 days in a year. Do not round intermediate calculations. Enter your answers as a percent rounded to 2 decimal places, e.g., 32.16.)

Stated Rate (APR) Number of Times Compounded Effective Rate (EAR)
8.1 % Quarterly %
17.1 Monthly
13.1 Daily
10.1 Infinite
Answer #1

EAR=(1+APR/m)^m-1
where m=compounding periods

1.EAR=(1+0.081/4)^4-1

=8.35%(Approx).

2.EAR=(1+0.171/12)^12-1

=18.51%(Approx).
3.EAR=(1+0.131/365)^365-1

=13.99%(Approx).

4.EAR=(e)^APR-1

where e=2.71828

EAR=(2.71828)^0.101-1

=10.63%(Approx).